OpenStudy (tylerbanks2001):

The heat in the house is set to keep the minimum and maximum temperatures (in degrees Fahrenheit) according to the equation |x – 72.5| = 4. What are the minimum and maximum temperatures in the house? 72.5°F and 76.5°F 68.5°F and 76.5°F 70.5°F and 74.5°F 72.5°F and 74.5°F

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. |x−72.5|=4 Solve Absolute Value. |x−72.5|=4 We know eitherx−72.5=4orx−72.5=−4 x−72.5=4(Possibility 1) x−72.5+72.5=4+72.5(Add 72.5 to both sides) x=76.5 x−72.5=−4(Possibility 2) x−72.5+72.5=−4+72.5(Add 72.5 to both sides) x=68.5 Answer: x=76.5 or x=68.5